A radio-frequency amplifier is designed to amplify signals in the frequency range 79.0 MHz to 121 MHz ▼ Part A What is the smallest duration radio-frequency pulse that can be amplified without distortion? Express your answer with the appropriate units Value Units Request Answer
Answer #1

Given,

f1 = 79 MHz ; f2 = 121 MHz

A)the smallest frequency will be;

fs = f2 - f1

fs = 121 - 79 = 42 MHz

Hence, fs = 42 MHz (Mega-Hz)